# Sampling Density for Gabor Phase Retrieval

### v2

@@ 1. Introduction and main results
The phase retrieval arises in recovering a signal from its intensity measurements, which is widely applied in X-ray crystallography, optical imaging, electron microscopy and many other fields [44]. Recently, phase retrieval has been studied in various aspects, which include phaseless sampling in shift invariant spaces [10, 11, 12, 13, 16, 20, 39, 40, 42, 45, 46], phase retrieval for quaternion signals [41, 52], affine linear measurements [17, 18, 32], masked Fourier measurements [38, 37] and sparse signals [50], and the stability of recovery [2, 3, 5, 9, 21, 27, 30, 31, 51, 55].
@@
For general window functions and bandlimited signals, Alaifari and Wellershoff [6] showed that an equally spaced sequence in certain line does phase retrieval for real-valued bandlimited functions. Zhang, Guo, Liu and Li [54] extended this result for vector-valued functions. Alaifari and Wellershoff [7] proved that the bandlimitedness assumption is necessary. Matthias [47] showed that two sequences taken respectively from two lines do Gabor phase retrieval for bandlimited functions. See also Grohs and Liehr [22] and Wellershoff [48] for the phase retrieval of complex-valued compactly supported functions, and [28, 29, 24, 1, 4] for the stability of Gabor phase retrieval.
@@
We see from the frame theory that for a well-chosen window function, any function in $L^{2}(\mathbb{R})$ is uniquely determined by its sampled values on a lattice with sufficiently large density [14, 15, 19]. However, things become quite different for phase retrieval. In [23], Grohs and Liehr showed that for any window function, any lattice does not do phase retrieval. Recently, Grohs, Liehr and Rathmair [26] gave a positive result which shows that for well chosen four window functions, a lattice with certain density does phase retrieval.
@@
Suppose that $\Lambda=\cup_{l=1}^{2}\big((t_{0},\omega_{0})+a\mathbb{Z}^{\nu}(\sin\theta_{l},\cos\theta_{l})\big)$, where $a$ and $\nu$ are positive number, $(t_{0},\omega_{0})\in\mathbb{R}^{2}$. If $(\theta_{1}-\theta_{2})/\pi\not\in\mathbb{Q}$ and either $\nu<1/2$ or $\nu=1/2$ with $a<1$, then $\Lambda$ does Gabor phase retrieval.
@@ 4.2. Phase retrieval with sampling sequences of uniform density
Suppose that $\Lambda=\cup_{l=1}^{3}\big\{(t_{l},\omega_{l})\pm\lambda_{n}(\sin\theta_{0},\cos\theta_{0}):\,n\geq 1\big\}$, where $\theta_{0}\in[0,2\pi]$, $(t_{1},\omega_{1})$, $(t_{2},\omega_{2})$ and $(t_{3},\omega_{3})$ are three points for which

### v3

@@ 1. Introduction and main results
Phase retrieval arises in recovering a signal from its intensity measurements, which is widely applied in X-ray crystallography, optical imaging, electron microscopy and many other fields [44]. Recently, phase retrieval has been studied in various aspects, which include phaseless sampling in shift invariant spaces [10, 11, 12, 13, 16, 20, 39, 40, 42, 45, 46], phase retrieval for quaternion signals [41, 52], affine linear measurements [17, 18, 32], masked Fourier measurements [38, 37] and sparse signals [50], and the stability of recovery [2, 3, 5, 9, 21, 27, 30, 31, 51, 55].
@@
For general window functions and bandlimited signals, Alaifari and Wellershoff [6] showed that an equally spaced sequence in certain line does phase retrieval for real-valued bandlimited functions. Zhang, Guo, Liu and Li [54] extended this result for vector-valued functions. Matthias [47] showed that two sequences taken respectively from two lines do Gabor phase retrieval for bandlimited functions. See also Grohs and Liehr [22] and Wellershoff [48] for the phase retrieval of complex-valued compactly supported functions, and [28, 29, 24, 1, 4] for the stability of Gabor phase retrieval.
@@
We see from the frame theory that for a well-chosen window function, any function in $L^{2}(\mathbb{R})$ is uniquely determined by its sampled values on a lattice with sufficiently large density [14, 15, 19]. However, things become quite different for phase retrieval. Alaifari and Wellershoff [7] proved that any lattice does not do Gabor phase retrieval. And in [23], Grohs and Liehr showed that for any window function, any lattice does not do phase retrieval. Recently, Grohs, Liehr and Rathmair [26] gave a positive result which shows that for well chosen four window functions, a lattice with certain density does phase retrieval.
@@
Suppose that $\Lambda=\cup_{l=1}^{2}\big((t_{0},\omega_{0})+a\mathbb{Z}^{\nu}(\sin\theta_{l},\cos\theta_{l})\big)$, where $a$ and $\nu$ are positive numbers, $(t_{0},\omega_{0})\in\mathbb{R}^{2}$. If $(\theta_{1}-\theta_{2})/\pi\not\in\mathbb{Q}$ and either $\nu<1/2$ or $\nu=1/2$ with $a<1$, then $\Lambda$ does Gabor phase retrieval.
@@ 4.2. Phase retrieval with sampling sequences of uniform density
Suppose that $\Lambda=\cup_{l=1}^{3}\big\{(t_{l},\omega_{l})\pm\lambda_{n,l}(\sin\theta_{0},\cos\theta_{0}):\,n\geq 1\big\}$, where $\theta_{0}\in[0,2\pi]$, $(t_{1},\omega_{1})$, $(t_{2},\omega_{2})$ and $(t_{3},\omega_{3})$ are three points for which